# Sim-Grasp: Learning 6-DOF Grasp Policies for Cluttered Environments Using a Synthetic Benchmark

### v1

@@ I INTRODUCTION
Robotic grasping is a fundamental problem in robotics research, focusing on the manipulation of objects with varying shapes and sizes in diverse environments. Grasping objects in cluttered environments [1] presents a complex challenge that requires the robotic system to accurately detect and localize unknown objects. As the demand for fully autonomous robotic systems capable of handling challenging tasks increases, traditional stationary industrial robotic arms may not be suitable for grasping objects in unstructured and dynamic environments. Consequently, the development of mobile manipulation platforms [2, 3] has gained significant attention due to their effectiveness in performing grasping tasks in such environments. Mobile grasping has become a hot topic in recent years [4], as researchers dedicate efforts to create a “GPT moment” for robots, empowering them with the ability to interact with the physical world. This paradigm shift is exemplified by Google DeepMind’s RT-2 [5], which highlights the immense potential of combining advanced language models with robotic systems, enabling them to understand and execute complex tasks in a more intuitive and human-like manner. Toyota has made significant strides in deploying mobile manipulators for real-world applications, such as grocery shopping [6]. These endeavors show the potential of robotic grasping not only in controlled settings but also in more challenging scenarios. The rapid advancements in artificial intelligence (AI) research have propelled intelligent robotic grasping to the forefront of the robotics community. Among the various grasping mechanisms, the two-finger gripper, also known as a parallel jaw gripper, has garnered significant attention due to its simplicity, robustness, and versatility. It enables robotic systems to handle a diverse array of objects commonly encountered in various tasks. In this work, we aim to explore the frontiers of grasping policies and create multi-modal grasping policies that can utilize text prompts or box prompts as input to perform open-set grasping. We propose Sim-Grasp system (Fig. 1), a deep learning-based system that utilizes a two-finger gripper to pick up novel objects from cluttered environments. It consists of three components: (1) Sim-Grasp-Dataset: a large-scale synthetic dataset for cluttered environments, (2) Sim-GraspNet: a 6 DOF grasp estimation network, and (3) Sim-Grasp-Policies: multi-modal grasping policies that include an object-agnostic picking mode, as well as text prompt and box prompt picking modes. The primary contributions of our work include:
@@ II-A Deep Learning for Grasping
Estimating 6-DOF grasp poses is crucial for robotic manipulation, especially in cluttered environments. 6-DOF grasp pose estimation has emerged as a more advanced and versatile approach compared to planar grasping. While planar grasping methods focus on predicting grasp poses in a 2D plane, typically using RGB or depth images, 6-DOF grasp pose estimation aims to predict the complete 6-DOF pose of the gripper, including its position and orientation in 3D space. By considering the full 6-DOF pose [11], robots can adapt to objects with complex geometries, asymmetric shapes, and various orientations. This is particularly important in cluttered environments, where objects may be partially occluded or positioned in unconventional ways. Planar grasping methods, on the other hand, are limited to objects that can be grasped from a top-down perspective and may struggle with more diverse object arrangements. Some methods [12] adopt a two-stage pipeline, where grasp candidates are first sampled and then evaluated using a learned model. Others [13, 14, 15] propose end-to-end architectures that directly predict grasp poses from input point clouds. While recent works in 6-DOF grasp pose estimation have achieved promising results, they often face limitations in terms of generalization and adaptability to diverse environments and objects. Many of these methods are constrained to specific settings, such as fixed vision systems, particular camera types, or a limited set of selected items. These constraints hinder their ability to perform open-set grasping, where the robot must handle novel objects and environments not seen during training.
@@ III Sim-Grasp-Dataset
We present a novel approach to creating a large-scale synthetic dataset Sim-Grasp-Dataset (Fig. 2) tailored for parallel jaw grippers, which builds upon our previous work, Sim-Suction [23]. The cluttered environments in Sim-Grasp-Dataset are identical to those in Sim-Suction, featuring the same 1,550 unique objects and 500 cluttered scenarios. To adapt these environments specifically for two-finger grasping modality, we introduce an approach-based sampling scheme coupled with dynamic evaluation. This process generates a comprehensive set of 7.8 million 6D grasping labels, enabling the development of robust grasping algorithms for parallel jaw grippers.
@@
To begin the sampling process (Fig. 3), we employ iterative Farthest Point Sampling (FPS) [24] to select representative points from each object’s point cloud. For each selected point, we compute the surface normal, which aids in defining the gripper’s approach direction. We have chosen sampling parameter $\alpha$ to be in $pi/6$ intervals, a configuration found to provide robust coverage according to findings in [25]. This setup establishes the approach direction $V$ for the gripper. Once the approach direction $V$ is established, and with the gripper’s z-axis aligned accordingly, we perform a collision check by casting rays along $V$ to detect any potential collisions. This process is applied only to collision-free approach directions. We then reconstruct the remaining two axes of the gripper and proceed to sample around $V$ by varying the angle $A$ and standoff distance $D$ to generate initial grasp candidates. After generating these candidates, we further refine them in the Isaac Sim simulator [26] environment, where the gripper is simplified using four box components representing the two fingers, the base, and the tail. We check the overlap of these boxes corresponding to the gripper pose with the object’s triangle mesh and employ a heuristic to ensure that the area close to the gripper’s fingers is free from obstructions. We annotate the overlap quality of each grasp candidate using the Intersection over Union (IOU) metric, which evaluates the overlap between the gripper boxes and the object mesh. After obtaining grasp candidates that successfully pass the overlap check, we further evaluate and label each grasp using Isaac Sim to ensure a realistic physics simulation. We mount the gripper on a 7-DoF UR10 robot to simulate the dynamic behavior accurately. Each object within the cluttered environment is modeled as a rigid body with distinct mass and physical properties, allowing for the simulation of realistic interactions among the objects. Unlike previous approaches [20, 9, 8], which derive cluttered scene labels from single object results through simple label mapping, our method directly annotates the cluttered environment. By taking into account the dynamics of each object’s interactions, we mitigate the risk of generating inaccurate ground truth labels, particularly in scenarios where objects are stacked or densely arranged. The effectiveness of this dynamic evaluation approach has been thoroughly discussed in our previous work, Sim-Suction [23]. By leveraging a high-fidelity physics simulation and directly annotating the cluttered environment, we provide a more accurate and realistic assessment of grasp efficacy in complex scenarios.
@@ IV-C Loss Function
$\mathcal{L}=\lambda_{\text{aff}}L_{\text{reg}}^{\text{aff}}+\lambda_{\text{dir}}\mathcal{L}_{\text{cls}}^{\text{dir}}+\lambda_{\text{score}}\mathcal{L}_{\text{cls}}^{\text{score}},$
@@ IV-D Multi-Modal Grasping Policies
Our Sim-Grasp-Policies are quite versatile and can handle a variety of tasks. We design the policies to work in three different modes: object-agnostic grasping, target picking with text prompts, and target picking with box prompts. In the object-agnostic grasping mode, the system can pick up a wide array of objects without the need for specific instructions, making it suitable for general-purpose robotic manipulation tasks. This modality is the direct output of Sim-GraspNet. For target picking with text prompts, we use a pre-trained Grounding Dino [29] module as a zero-shot object detector that takes text input and generates object bounding boxes. These bounding boxes are then processed into a pre-trained zero-shot Segment Anything (SAM) module [30], which generates valid segmentation masks for the specified objects. This modality allows users to specify target objects using natural language descriptions. It is particularly useful in scenarios that require picking specific objects based on their semantic attributes. In the target picking with box prompts mode, users can directly provide bounding box prompts into the SAM module to generate segmentation masks for the objects of interest. This modality offers a high degree of control over the grasping policy, especially for items that are difficult to describe using natural language. The segmentation masks produced by the SAM module, along with the 6-DOF grasp poses generated by the Sim-GraspNet backbone for the M seeds, are merged in the grasp refinement module. This module ensures that each grasp candidate is object-aware and incorporates a collision detector submodule to avoid potential collisions during grasping. Furthermore, a non-maximum suppression (NMS) submodule is used to merge grasp candidates that are in close proximity to each other for the same object, thereby ensuring a diverse coverage of the entire scene when outputting the top percentage of candidates.
@@ V-A Online Ablation study
To establish a baseline for comparison, we employ the Grasp Pose Detection Network (GPD) [7] to detect 6-DOF grasp poses. We then evaluate the performance of Sim-GraspNet trained on two different types of point cloud data: single-view (SV) and multi-view merged (MV). In our dataset, we have 500 cluttered environments. The MV dataset consists of 500 point clouds, each created by merging multiple views of a single environment. On the other hand, the SV dataset contains 100,000 point clouds, generated by capturing 200 unique views for each environment, resulting in a total of 100,000 point clouds with different camera viewpoints. Both MV and SV point clouds are augmented in the training process by random scaling and rotation. The performance is assessed in the Isaac Sim Simulator using Average Precision (AP), which is calculated as the ratio of successful grasps to total grasp candidates. Table II presents the results, where similar objects refer to those sharing common characteristics with the training dataset but differing in scale, while novel objects are those introduced during the testing phase and not included in the training dataset. Both SV and MV versions of Sim-GraspNet outperform the baseline, highlighting our approach’s effectiveness. Interestingly, Sim-GraspNet trained on SV data slightly outperforms the MV-trained version, contrasting with our findings in Sim-Suction [23], where MV training yielded better results. This difference can be attributed to the distinct nature of suction and grasping. Suction relies on the perceived area, and training on MV point clouds enhances the model’s understanding of continuous surfaces due to less occlusion. On the other hand, Grasping involves more complex object geometry interactions. When inferring grasps on a single-view point cloud, the model lacks information about the unseen half of the object. MV training may result in predicting grasps on object edges, which can be valid or invalid depending on the hidden geometry. In contrast, SV training encourages the model to make inferences based on the available partial point cloud, resulting in more reliable grasp predictions.
@@ VI CONCLUSIONS & FUTURE WORK
In this paper, we present Sim-Grasp, a deep learning-based two-finger grasping system for objects in cluttered environments. Experiments conducted on a mobile manipulation platform demonstrate that Sim-Grasp, learned from the synthetic dataset Sim-Grasp-Dataset, achieves a robust success rate in real-world cluttered environments with dynamic viewpoints. In the future, we intend to incorporate closed-loop feedback mechanisms to further improve the system performance. This will involve the integration of tactile sensors and load cells to provide real-time data on grasp success, object weight, and manipulation dynamics.

### v2

@@ Abstract
Index Terms:
Grasping, Mobile Manipulation, Deep Learning in Grasping and Manipulation, Data Sets for Robot Learning
@@ I INTRODUCTION
Robotic grasping is a fundamental problem in robotics research, focusing on the manipulation of objects with varying shapes and sizes in diverse environments. Grasping objects in cluttered environments [1] presents a complex challenge that requires the robotic system to accurately detect and localize unknown objects. As the demand for fully autonomous robotic systems capable of handling challenging tasks increases, traditional stationary industrial robotic arms may not be suitable for grasping objects in unstructured and dynamic environments. Consequently, the development of mobile manipulation platforms [2, 3, 4] has gained significant attention due to their effectiveness in performing grasping tasks in such environments. The rapid advancements in artificial intelligence (AI) research have propelled intelligent robotic grasping to the forefront of the robotics community. Among the various grasping mechanisms, the two-finger gripper, also known as a parallel jaw gripper, has garnered significant attention due to its simplicity, robustness, and versatility. It enables robotic systems to handle a diverse array of objects commonly encountered in various tasks. In this work, we aim to explore the frontiers of grasping policies and create multi-modal grasping policies that can utilize text prompts or box prompts as input to perform open-set grasping. We propose Sim-Grasp system (Fig. 1), a deep learning-based system that utilizes a two-finger gripper to pick up novel objects from cluttered environments. It consists of three components: (1) Sim-Grasp-Dataset: a large-scale synthetic dataset for cluttered environments, (2) Sim-GraspNet: a 6 DOF grasp estimation network, and (3) Sim-Grasp-Policies: multi-modal grasping policies that include an object-agnostic picking mode, as well as text prompt and box prompt picking modes. The primary contributions of our work include:
@@ II-A Deep Learning for Grasping
Estimating 6-DOF grasp poses is crucial for robotic manipulation, especially in cluttered environments. 6-DOF grasp pose estimation has emerged as a more advanced and versatile approach compared to planar grasping. While planar grasping methods focus on predicting grasp poses in a 2D plane, typically using RGB or depth images, 6-DOF grasp pose estimation aims to predict the complete 6-DOF pose of the gripper, including its position and orientation in 3D space. By considering the full 6-DOF pose [9, 10], robots can adapt to objects with complex geometries, asymmetric shapes, and various orientations. This is particularly important in cluttered environments, where objects may be partially occluded or positioned in unconventional ways. Planar grasping methods, on the other hand, are limited to objects that can be grasped from a top-down perspective and may struggle with more diverse object arrangements. Some methods [11] adopt a two-stage pipeline, where grasp candidates are first sampled and then evaluated using a learned model. Others [12, 13, 14] propose end-to-end architectures that directly predict grasp poses from input point clouds. While recent works in 6-DOF grasp pose estimation have achieved promising results, they often face limitations in terms of generalization and adaptability to diverse environments and objects. Many of these methods are constrained to specific settings, such as fixed vision systems, particular camera types, or a limited set of selected items. These constraints hinder their ability to perform open-set grasping, where the robot must handle novel objects and environments not seen during training.
@@ III Sim-Grasp-Dataset
We present a novel approach to creating a large-scale synthetic dataset Sim-Grasp-Dataset (Fig. 2) tailored for parallel jaw grippers, which builds upon our previous work, Sim-Suction [22]. The cluttered environments in Sim-Grasp-Dataset are similar to those in Sim-Suction, featuring 1,550 unique objects and 500 cluttered scenarios. The object models come from ShapeNet [23], Google Scanned objects [24], NVIDIA Omniverse Assets [25], open source 3D models, and Adversarial Objects in DexNet [6]. To adapt these environments specifically for two-finger grasping modality, we introduce an approach-based sampling scheme coupled with dynamic evaluation. This process generates a comprehensive set of 7.8 million 6D grasping labels, enabling the development of robust grasping algorithms for parallel jaw grippers.
@@
To begin the sampling process (Fig. 3), we employ iterative Farthest Point Sampling (FPS) [26] to select representative points from each object’s point cloud. For each selected point, we compute the surface normal, which aids in defining the gripper’s approach direction. We have chosen sampling parameter $\alpha$ to be in $pi/6$ intervals, a configuration found to provide robust coverage according to findings in [27]. This setup establishes the approach direction $V$ for the gripper. Once the approach direction $V$ is established, and with the gripper’s z-axis aligned accordingly, we perform a collision check by casting rays along $V$ to detect any potential collisions. This process is applied only to collision-free approach directions. We then reconstruct the remaining two axes of the gripper and proceed to sample around $V$ by varying the angle $A$ and standoff distance $D$ to generate initial grasp candidates. After generating these candidates, we further refine them in the Isaac Sim simulator [28] environment, where the gripper is simplified using four box components representing the two fingers, the base, and the tail. We check the overlap of these boxes corresponding to the gripper pose with the object’s triangle mesh and employ a heuristic to ensure that the area close to the gripper’s fingers is free from obstructions. We annotate the overlap quality of each grasp candidate using the Intersection over Union (IOU) metric, which evaluates the overlap between the gripper boxes and the object mesh. After obtaining grasp candidates that successfully pass the overlap check, we further evaluate and label each grasp using Isaac Sim to ensure a realistic physics simulation. We mount the gripper on a 7-DoF UR10 robot, modeling each object as a rigid body with unique properties. Unlike previous works [19, 7, 6] that use simple label mapping, we directly annotate cluttered environments by considering object dynamics. The physics engine simulates essential parameters like velocity, acceleration, and mass, accurately modeling complex interactions in cluttered scenes. This approach captures realistic scenarios, such as failed grasps due to dynamic interactions when attempting to grasp objects from the bottom of a pile. Our dataset generation process mimics real-world conditions, producing high-quality labels that enable effective transfer to real-world scenarios. This dynamic evaluation approach, validated in our previous work Sim-Suction [22], provides a more accurate assessment of grasp efficacy in complex scenarios compared to simple label mapping methods.
@@ IV-C Loss Function
$\mathcal{L}=\lambda_{\text{aff}}\mathcal{L}_{\text{reg}}^{\text{aff}}+\lambda_{\text{dir}}\mathcal{L}_{\text{cls}}^{\text{dir}}+\lambda_{\text{score}}\mathcal{L}_{\text{cls}}^{\text{score}}$
@@ IV-D Multi-Modal Grasping Policies
Our Sim-Grasp-Policies are quite versatile and can handle a variety of tasks. We design the policies to work in three different modes: object-agnostic grasping, target picking with text prompts, and target picking with box prompts. In the object-agnostic grasping mode, the system can pick up a wide array of objects without the need for specific instructions, making it suitable for general-purpose robotic manipulation tasks. This modality is the direct output of Sim-GraspNet. For target picking with text prompts, we use a pre-trained Grounding Dino [31] module as a zero-shot object detector that takes text input and generates object bounding boxes. These bounding boxes are then processed into a pre-trained zero-shot Segment Anything (SAM) module [32], which generates valid segmentation masks for the specified objects. This modality allows users to specify target objects using natural language descriptions. It is particularly useful in scenarios that require picking specific objects based on their semantic attributes. In the target picking with box prompts mode, users can directly provide bounding box prompts into the SAM module to generate segmentation masks for the objects of interest. This modality offers a high degree of control over the grasping policy, especially for items that are difficult to describe using natural language. The segmentation masks produced by the SAM module, along with the 6-DOF grasp poses generated by the Sim-GraspNet backbone for the M seeds, are merged in the grasp refinement module. This module ensures that each grasp candidate is object-aware and incorporates a point-cloud based model-free collision submodule to filter out potential collisions during grasping. Furthermore, a non-maximum suppression (NMS) submodule is used to merge grasp candidates that are in close proximity to each other for the same object, thereby ensuring a diverse coverage of the entire scene when outputting the top percentage of candidates.
@@ V-A Online Ablation study
To establish a baseline for comparison, we employ the Grasp Pose Detection Network (GPD) [5] to detect 6-DOF grasp poses. We then evaluate the performance of Sim-GraspNet in a single-view setting trained on two different types of point cloud data: single-view (SV) and multi-view merged (MV). In our dataset, we have 500 cluttered environments. The MV dataset consists of 500 point clouds, each created by merging multiple views of a single environment. On the other hand, the SV dataset contains 100,000 point clouds, generated by capturing 200 unique views for each environment, resulting in a total of 100,000 point clouds with different camera viewpoints. Both MV and SV point clouds are augmented in the training process by random scaling and rotation. The performance is assessed in the Isaac Sim Simulator using Average Precision (AP), which is calculated as the ratio of successful grasps to total grasp candidates. Table II presents the results, where similar objects refer to those sharing common characteristics with the training dataset but differing in scale, while novel objects are those introduced during the testing phase and not included in the training dataset. Both SV and MV versions of Sim-GraspNet outperform the baseline, highlighting our approach’s effectiveness. Interestingly, Sim-GraspNet trained on SV data slightly outperforms the MV-trained version, contrasting with our findings in Sim-Suction [22], where MV training yielded better results. This difference can be attributed to the distinct nature of suction and grasping. Suction relies on the perceived area, and training on MV point clouds enhances the model’s understanding of continuous surfaces due to less occlusion. On the other hand, Grasping involves more complex object geometry interactions. When inferring grasps on a single-view point cloud, the model lacks information about the unseen half of the object. MV training may result in predicting grasps on object edges, which can be valid or invalid depending on the hidden geometry. In contrast, SV training encourages the model to make inferences based on the available partial point cloud, resulting in more reliable grasp predictions.
@@ VI CONCLUSIONS & FUTURE WORK
In this paper, we present Sim-Grasp, a deep learning-based two-finger grasping system for objects in cluttered environments. Experiments conducted on a mobile manipulation platform demonstrate that Sim-Grasp, learned from the synthetic dataset Sim-Grasp-Dataset, achieves a robust success rate in real-world cluttered environments with dynamic viewpoints. Our system uses point cloud data to generate 6D grasping poses, effective for objects with well-defined geometries and opaque surfaces. However, it struggles with transparent objects due to insufficient point cloud data from the Fetch head camera. For deformable objects, the lack of force feedback sensors (load cell or tactile) poses challenges in handling delicate items, risking damage during manipulation. Nevertheless, the system can grasp softer deformable objects like plush toys relatively well. In the future, we intend to incorporate closed-loop feedback mechanisms to further improve the system performance. This will involve the integration of tactile sensors and load cells to provide real-time data on grasp success, object weight, and manipulation dynamics.